# Orak: A Foundational Benchmark for Training and Evaluating LLM Agents on Diverse Video Games

### v1

@@ 4 Fine-tuning: Aligning Pre-trained LLMs into Game Agents
$X^{\text{ref}}\!\!$ Analyze Mario’s past action using state difference and provide critiques for improving his action. You should only respond in the format as below: ######\#\#\## # # Self-reflection (Describe self-reflection here) $X^{\text{act}}$ (Retrieve the self-reflection from memory) Analyze the current game state, and decide the best action. You should only respond in the format as below: ######\#\#\## # # Action Jump Level: n (where n is an integer from 0 to 6)
$S$#####\#\#\## # # Past Game State Mario at (100,100), Bricks at (120, 100), (120, 150) ######\#\#\## # # Current Game State Mario at (100,100), Bricks at (120, 100), (120, 150) $S$#####\#\#\## # # Game State Mario at (100,100) Position of all objects: - Bricks at (120, 100), (120, 150) $\ldots$
$Y^{\text{ref}}$#####\#\#\## # # Self-reflection Mario is blocked by bricks. Jump higher to get past.  $Y^{\text{act}}$#####\#\#\## # # Action Jump Level: 6.
@@
We collect the fine-tuning dataset from expert LLMs, e.g., GPT-4o and o3-mini, playing 12 games in Orak using several agentic modules. This dataset with environment interaction trajectories encapsulates meta-knowledge on how to use the agentic strategies to solve diverse game genres.
[…]
Data Selection. For each game in Orak, we first collect $N$ gameplay trajectories $\mathcal{T}$, where $N$ varies depending on the game. To ensure high-quality data, we select the trajectories with high game scores until the number of LLM inferences exceeds 900. All selected trajectories follow the ‘reflection-planning-action’ sequence, so that we have around 300 samples for each agent module. By performing data selection on all 12 games, our fine-tuning set consists of approximately 10k samples.
Data Augmentation. To enhance the linguistic diversity, we augment each data sample $\tau$ by paraphrasing. We prompt GPT-4o to rephrase the game prompt $X^{a}$ while preserving all game-related information, generating 10 augmented samples for each sample $\tau$. See Appendix O for details.
@@ 5.1 Experiment Setup
Models. On all games in Orak, we validate the performance of 12 LLMs using states provided in text format. The models include 6 open-source LLMs: LLaMA-3.2-1B/3B [66], Qwen-2.5-3B/7B [67], and Minitron-4B/8B [68], and 6 proprietary LLMs: GPT-4o-mini, GPT-4o [69], o3-mini, Gemini-2.5-pro [70], Claude-3.7-sonnet [71], and DeepSeek-R1 [72]. In addition, we study the effects of incorporating image inputs—either alongside text or as the sole input—on multi-modal LLMs: GPT-4o, Gemini-2.5-pro, and Claude-3.7.
Default Agentic Strategies. For each game in Orak, we select the most effective agentic strategy over GPT-4o and o3-mini as the default agent strategy. Specifically, For Street Fighter III, HerStory, Darkest Dungeon, and 2048, we use a ‘zero-shot’ action inference agent. For Super Mario, Pokémon Red, Stardew Valley, StarCraft II, Slay the Spire, and Baba Is You, we use a ‘reflection-planning’ agent that sequentially performs self-reflection, subtask planning, and action inference, integrated with memory at each game step. For AceAttorney, we use a ‘reflection’ agent. For Minecraft, we use a ‘skill-management’ agent that further includes knowledge retrieval and skill management in the reflection-planning-action agent, following [20].
Metrics and Implementation Details. For each game, we report the normalization score rather than the absolute score, i.e., the game score is normalized by the maximum game score. We report the average score of 3 to 20 trials for each game. More detailed metrics and LLM hyperparameter configurations are in Appendix P.
@@ 5.2 LLM Gameplay Performance
Llama-3.2-1B 0.0$\pm$0.0 18.7$\pm$8.6 1.3$\pm$2.2 2.1$\pm$1.2 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 6.7$\pm$11.5 0.0$\pm$0.1 10.8
Llama-3.2-3B 13.3$\pm$5.8 31.8$\pm$10.1 4.6$\pm$1.3 4.2$\pm$1.1 0.0$\pm$0.0 47.5$\pm$39.2 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 20.0$\pm$0.0 0.3$\pm$0.2 8.9
Qwen-2.5-3B 20.0$\pm$0.0 23.4$\pm$14.1 20.0$\pm$17.4 1.2$\pm$1.1 0.0$\pm$0.0 44.8$\pm$22.2 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 13.3$\pm$11.5 0.1$\pm$0.1 9.5
Qwen-2.5-7B 16.7$\pm$11.5 27.2$\pm$9.6 9.3$\pm$0.2 8.5$\pm$1.9 0.0$\pm$0.0 88.8$\pm$2.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 5.0$\pm$0.0 20.0$\pm$0.0 0.6$\pm$0.4 8.1
Minitron-4B 16.7$\pm$11.5 24.4$\pm$6.0 35.7$\pm$4.5 4.5$\pm$2.2 0.0$\pm$0.0 0.0$\pm$0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 20.0$\pm$0.0 0.1$\pm$0.0 9.0
Minitron-8B 23.3$\pm$5.8 31.3$\pm$12.8 29.9$\pm$3.6 8.2$\pm$1.8 0.0$\pm$0.0 63.8$\pm$30.4 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 20.0$\pm$0.0 0.7$\pm$0.7 7.7
GPT-4o-mini 16.7$\pm$11.5 28.8$\pm$8.8 28.4$\pm$2.8 21.1$\pm$5.5 0.0$\pm$0.0 81.3$\pm$5.8 46.0$\pm$7.0 18.9$\pm$32.7 75.0$\pm$50.0 3.3$\pm$2.9 13.3$\pm$11.5 1.1$\pm$1.0 6.8
GPT-4o 29.7$\pm$14.3 34.1$\pm$14.2 85.3$\pm$1.5 64.0$\pm$5.1 38.9$\pm$9.6 93.4$\pm$1.5 71.0$\pm$7.0 95.7$\pm$5.7 100.0$\pm$0.0 23.6$\pm$22.1 20.0$\pm$0.0 5.6$\pm$1.5 2.9
o3-mini 33.3$\pm$15.3 34.9$\pm$14.6 91.7$\pm$1.5 66.3$\pm$3.6 0.0$\pm$0.0 89.0$\pm$2.1 75.0$\pm$0.0 64.7$\pm$18.8 25.0$\pm$50.0 15.0$\pm$0.0 73.3$\pm$46.2 25.3$\pm$7.3 3.3
Gemini-2.5-pro 13.3$\pm$11.5 38.0$\pm$14.4 55.7$\pm$3.4 67.3$\pm$3.3 83.3$\pm$0.0 93.7$\pm$1.6 75.0$\pm$0.0 69.6$\pm$11.9 100.0$\pm$0.0 51.9$\pm$31.9 73.3$\pm$46.2 5.1$\pm$2.5 2.8
Claude-3.7 16.7$\pm$11.5 31.7$\pm$8.2 81.9$\pm$1.6 62.6$\pm$2.6 63.9$\pm$19.2 89.9$\pm$2.5 75.0$\pm$0.0 63.0$\pm$24.6 50.0$\pm$57.7 15.0$\pm$0.0 46.7$\pm$46.2 5.3$\pm$2.7 4.2
Deepseek-R1 20.0$\pm$0.0 28.7$\pm$13.2 83.3$\pm$1.5 66.9$\pm$3.9 75.0$\pm$0.0 91.7$\pm$1.1 41.7$\pm$0.0 77.7$\pm$13.6 50.0$\pm$57.7 24.9$\pm$17.1 20.0$\pm$0.0 11.5$\pm$3.4 4.0
Table 3: Performance of LLMs on Orak with default agentic strategies. The best scores for each game are highlighted in bold, and the average ranking for each LLM across all games is reported.
Table 3 shows the gameplay performance of LLMs on Orak with default agentic strategies. Overall, proprietary LLMs outperform open-source LLMs across all games in most cases. Gemini-2.5-pro performs the best on average, ranking first in 8 out of 12 games, with the best average ranking of 2.7. o3-mini shows the second-best performance with the average ranking of 2.7, while showing the best in puzzle games, i.e., Baba Is You and 2048, which require strong mathematical and logical reasoning, and spatial understanding abilities. Most open-source LLMs show almost zero score on complex games such as Pokémon-red, Minecraft, Stardew Valley, StarCraft II, and Slay the Spire. Among open-source LLMs, Minitron-8B achieves the best average ranking of 7.7, and Qwen-2.5-7B shows the second-best average ranking of 8.1. See Appendices C-N for more detailed results of each game.
@@ 5.3 LLM Arena
Among 12 games in Orak, Street Fighter III and StarCraft II support two-player competitive modes. For Street Fighter III, we conduct pairwise battles among 8 LLMs with ‘zero-shot’ agent. Each pair competed in three rounds, and the agent winning 2 out of 3 rounds was declared the winner. To ensure a fair comparison, both agents were assigned the same character, Ken, in all matches. Figure 4(a) shows the relative win rates and Elo ratings. Interestingly, as opposed to the result in Section 5.2, Minitron-8B consistently outperforms all other LLMs and achieves the best Elo rating. This may imply that when multiple agents are involved in the environment, adversarial actions can significantly change the game dynamics. For StarCraft II, we conduct pairwise battles among 7 LLMs, with each pair competing in a single round. Both agents were assigned the same race, Protoss, in all matches. As shown in Figure 4(b), Claude-3.7-Sonnet performs the best with the highest Elo rating, while GPT-4o and Minitron-8B form the second group.
@@ 5.4 Ablation Study for Agentic Modules
LLaMA-3B Zeroshot 13.3$\pm$5.8 21.2$\pm$8.2 5.7$\pm$3.2 4.2$\pm$1.1 0.0$\pm$0.0 47.5$\pm$39.2 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 20.0$\pm$0.0 0.3$\pm$0.2 6.4
Reflection 30.0$\pm$17.3 32.4$\pm$8.6 4.6$\pm$1.3 4.4$\pm$1.3 0.0$\pm$0.0 47.3$\pm$39.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 20.0$\pm$0.0 0.0$\pm$0.0 5.6
Planning 20.0$\pm$0.0 27.0$\pm$8.4 4.6$\pm$1.3 5.1$\pm$1.0 0.0$\pm$0.0 56.3$\pm$23.6 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 20.0$\pm$0.0 0.1$\pm$0.1 6.2
Ref-Plan 16.7$\pm$20.8 31.8$\pm$10.1 3.8$\pm$0.0 5.4$\pm$0.4 0.0$\pm$0.0 57.0$\pm$31.6 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 20.0$\pm$0.0 0.1$\pm$0.2 6.1
GPT-4o Zeroshot 29.7$\pm$14.3 29.6$\pm$9.2 49.9$\pm$1.3 64.0$\pm$5.1 33.3$\pm$0.0 93.4$\pm$1.5 0.0$\pm$0.0 40.5$\pm$35.2 50.0$\pm$57.7 24.7$\pm$9.2 20.0$\pm$0.0 5.6$\pm$1.5 3.3
Reflection 23.3$\pm$20.8 32.3$\pm$15.5 85.3$\pm$1.5 61.3$\pm$0.4 36.1$\pm$4.8 85.2$\pm$10.3 50.0$\pm$0.0 18.3$\pm$5.2 0.0$\pm$0.0 41.3$\pm$19.6 20.0$\pm$0.0 3.5$\pm$2.9 3.0
Planning 30.0$\pm$26.5 29.4$\pm$12.5 52.7$\pm$0.5 59.2$\pm$5.7 33.3$\pm$0.0 82.0$\pm$8.6 13.0$\pm$0.0 64.6$\pm$23.7 50.0$\pm$57.7 35.3$\pm$9.2 20.0$\pm$0.0 6.0$\pm$5.5 3.1
Ref-Plan 23.3$\pm$20.8 34.1$\pm$14.2 52.8$\pm$0.5 61.8$\pm$4.9 38.9$\pm$9.6 91.6$\pm$2.5 50.0$\pm$0.0 95.7$\pm$5.7 100.0$\pm$0.0 36.0$\pm$25.5 20.0$\pm$0.0 7.0$\pm$5.7 2.2
@@ 5.5 Effect of Visual Input
We studied the visual input effect on GPT-4o, Gemini-2.5-pro, and Claude-3.7, as suitable open-source multimodal models were unavailable during experiments. Games were divided into two categories: Group 1 (Table 6) includes games where the provided textual game state can be derived from a single visual screenshot; for these, we evaluated Text-only, Image-only, and Both (text and image) input modalities. Group 2 (Table 6) comprises games where the textual state contains information beyond what is visible in the current frame (e.g., abstracted inventory lists or off-screen character/item details); for this group, we evaluated Text-only and Both inputs. Minecraft was excluded from this vision-based analysis due to challenges in programmatic screenshot capture within our setup.
As shown in Table 6, relying solely on Image-only input led to a substantial drop in performance. This was consistently reflected across all models, with average ranks deteriorating significantly compared to Text-only input. In contrast, as shown in Table 66, utilizing Both text and visual inputs produced mixed effects on game scores and model ranks. For instance, in games like Street Fighter III (Group 1), since on-screen visual details are challenging to fully convey textually, adding visual context significantly benefited Claude, increasing its score by $16.6$. Conversely, in narrative-heavy games such as Ace Attorney (Group 2), the same approach often proved detrimental; GPT-4o’s score, for example, dropped by $31.8$, and its average rank in that group fell from 2.9 (Text-only) to 4.7. This highlights that the impact of combining modalities varied considerably, with some scenarios showing improved ranks or scores while others demonstrated a decline.
@@
GPT-4o Text 29.7$\pm$14.3 34.1$\pm$14.1 95.7$\pm$5.7 20.0$\pm$0.0 5.6$\pm$1.5 3.0
Image 23.7$\pm$15.9 27.1$\pm$13.7 0.0 $\pm$0.0 6.7$\pm$11.5 1.8$\pm$1.1 7.4
Both 24.3$\pm$14.5 27.1$\pm$10.2 49.2$\pm$26.0 20.0$\pm$0.0 5.4$\pm$4.5 5.3
Gemini Text 13.3$\pm$11.5 38.0$\pm$13.4 69.6$\pm$11.9 73.3$\pm$46.2 5.1$\pm$2.5 4.6
Image 16.7$\pm$11.5 28.5$\pm$10.7 8.9$\pm$10.5 20.0$\pm$0.0 5.5$\pm$2.4 5.9
Both 20.0$\pm$10.0 40.9$\pm$9.6 70.6$\pm$7.1 86.7$\pm$23.1 3.1$\pm$2.6 3.6
Claude Text 16.7$\pm$11.5 28.7$\pm$13.2 63.0$\pm$24.6 46.7$\pm$46.2 5.3$\pm$2.7 4.9
Image 23.3$\pm$11.5 25.6$\pm$6.4 0.0$\pm$0.0 20.0$\pm$0.0 8.4$\pm$4.0 5.7
Both 33.3$\pm$5.8 22.6$\pm$6.3 58.5$\pm$1.2 20.0$\pm$0.0 6.7$\pm$0.9 4.6
Table 5: Comparison across modality (Group 1).
@@
GPT-4o Text 85.3$\pm$1.5 64.0$\pm$5.1 38.9$\pm$9.6 93.4$\pm$1.5 100.0$\pm$0.0 23.6$\pm$22.1 2.9
Both 53.5$\pm$1.7 40.4$\pm$29.4 41.7$\pm$8.3 92.2$\pm$3.0 50.0$\pm$57.7 23.6$\pm$22.1 4.7
Gemini Text 55.7$\pm$3.4 67.3$\pm$3.3 83.3$\pm$0.0 93.7$\pm$1.6 100.0$\pm$0.0 51.9$\pm$31.9 1.8
Both 52.6$\pm$0.8 64.7$\pm$2.4 83.3$\pm$0.0 92.2$\pm$1.8 100.0$\pm$0.0 26.2$\pm$19.4 2.8
Claude Text 81.9$\pm$1.6 62.6$\pm$2.6 63.9$\pm$19.2 89.9$\pm$2.5 50.0$\pm$57.7 15.0$\pm$0.0 4.5
Both 71.3$\pm$17.3 63.4$\pm$3.1 72.2$\pm$4.8 90.1$\pm$5.7 50.0$\pm$57.7 9.7$\pm$4.6 4.3
Table 6: Comparison across modality (Group 2).

### v2

@@ 4 Fine-tuning: Aligning Pre-trained LLMs into Game Agents
$X^{\text{ref}}\!\!$
Analyze Mario’s past action using state difference and
provide critiques for improving his action.
You should only respond in the format as below:
###\#\#\# Self-reflection
(Describe self-reflection here)
$X^{\text{act}}$
(Retrieve the self-reflection from memory)
Analyze the current game state, and decide the best action.
You should only respond in the format as below:
###\#\#\# Action
Jump Level: n (where n is an integer from 0 to 6)
$S$
###\#\#\# Past Game State
Mario at (100,100), Bricks at (120, 100), (120, 150)
###\#\#\# Current Game State
Mario at (100,100), Bricks at (120, 100), (120, 150)
$S$
###\#\#\# Game State
Mario at (100,100)
Position of all objects:
- Bricks at (120, 100), (120, 150) $\ldots$
$Y^{\text{ref}}$
###\#\#\# Self-reflection
Mario is blocked by bricks. Jump higher to get past.
$Y^{\text{act}}$
###\#\#\# Action
Jump Level: 6.
@@
We collect the fine-tuning dataset from expert LLMs, e.g., GPT-4o and o3-mini, playing all 12 games in Orak using several agentic modules. This dataset with environment interaction trajectories encapsulates meta-knowledge on how to use the agentic strategies to solve diverse game genres.
Data Format. We denote LLMs’ gameplay trajectory as $\mathcal{T}=\{\tau_{1},\ldots,\tau_{T}\}$, where $T$ is the number of game steps, and $\tau_{t}$ denotes the sequence of LLM inferences executed by agentic strategies at game step $t$. Each LLM inference sequence $\tau$ is represented as $\tau=\{(X^{a_{i}},S,Y^{a_{i}})\}_{i=1}^{n}$, where $a_{i}\in\{\text{`reflection'},\text{`planning'},\ldots,\text{`action'}\}$ is the $i$-th agentic module in sequence, $X^{a}$ is the prompt for agentic module $a$, $S$ is the game state, and $Y^{a}$ is the corresponding response of LLM. Table 2 shows detailed data examples of $\tau$.
Data Selection. For each game in Orak, we collect gameplay trajectories of expert LLMs, until we have more than 1000 LLM inference sequence $\tau$ in all $\mathcal{T}$ collected. To ensure high-quality data, we sort the collected $\mathcal{T}$ in terms of the game score and select the trajectories with the highest game scores until the number of selected $\tau$ exceeds 300. All selected trajectories follow the ‘reflection-planning-action’ sequence, so that we have around 300 samples for each agent module. By performing data selection on all 12 games, our fine-tuning set consists of approximately 11k samples.
Data Augmentation. To enhance the linguistic diversity, we augment each data sample $\tau$ by paraphrasing. We prompt GPT-4o to rephrase the game prompt $X^{a}$ while preserving all game-related information, generating 10 augmented samples for each sample $\tau$. See Appendix O for prompting details and the effect of the number of augmentations.
@@ 5.1 Experiment Setup
Models.
We validate the performance of 15 LLMs using states provided in text format. The models include 8 open-source LLMs: LLaMA-3.2-1B/3B (Grattafiori et al., 2024), LLaMA-3.3-72B (Grattafiori et al., 2024), Qwen-2.5-3B/7B/72B (Yang et al., 2024), and Minitron-4B/8B (Sreenivas et al., 2024), and 7 proprietary LLMs: GPT-5/4o/4o-mini (Achiam et al., 2023), o3-mini, Gemini-2.5-pro (Team et al., 2023), Claude-3.7-sonnet (Anthropic, 2025), and DeepSeek-R1 (Guo et al., 2025). In addition, we study the effects of incorporating image inputs on 5 multi-modal LLMs: Qwen2.5-vl-7B/32B (Bai et al., 2025), GPT-4o, Gemini-2.5-pro, and Claude-3.7-sonnet.
Default Agentic Strategies.
For each game in Orak, we select the most effective agentic strategy over GPT-4o as the default agent strategy. Specifically, For Street Fighter III, HerStory, Darkest Dungeon, and 2048, we use a ‘zero-shot’ action inference agent. For Super Mario, Pokémon Red, Stardew Valley, StarCraft II, Slay the Spire, and Baba Is You, we use a ‘reflection-planning’ agent that sequentially performs self-reflection, subtask planning, and action inference, integrated with memory at each game step. For AceAttorney, we use a ‘reflection’ agent. For Minecraft, we use a ‘skill-management’ agent that further includes knowledge retrieval and skill management in the reflection-planning-action agent, following (Wang et al., 2023).
Metrics and Implementation Details.
For each game, we report the normalization score rather than the absolute score, i.e., the game score is normalized by the maximum game score. We report the average score of 3 to 20 trials for each game. We also report the average score of 3 human novices. More detailed metrics and LLM hyperparameter configurations can be found in Appendices C-N.
@@ 5.2 LLM Gameplay Performance
[…]
Table 3: Performance of LLMs on Orak with default agentic strategies. The best scores for each game are highlighted in bold, and the average LLM rankings across all games are reported.
Table 3 shows the gameplay performance of LLMs on Orak with default agentic strategies. Overall, proprietary LLMs outperform open-source LLMs across all games in most cases. Gemini-2.5-pro performs the best on average, ranking first in 5 out of 12 games, with the best average ranking of 3.5. gpt-5/4o forms the second-best group with an average ranking of 3.6. GPT-5 shows superiority in puzzle games, i.e., Baba Is You and 2048, which require strong mathematical and logical reasoning, and spatial understanding abilities. Most small open-source LLMs, size under 8B, show almost zero score on complex games, i.e., Pokémon-red, Minecraft, Stardew Valley, StarCraft II, and Slay the Spire. Mid-sized open-source models, Llama3.3-70B and Qwen2.5-72B, show advances in moderately hard games, such as Stardew Valley and Slay the Spire, but their performance remains far below that of proprietary LLMs. See Appendices C-N for more detailed results on each game.
@@ 5.3 LLM Arena
Among 12 games in Orak, Street Fighter III and StarCraft II support two-player competitive modes. For Street Fighter III, we conduct pairwise battles among 8 LLMs with ‘zero-shot’ agent. Each pair competed in three rounds, and the agent winning 2 out of 3 rounds was declared the winner. To ensure a fair comparison, both agents were assigned the same character, Ken, in all matches. Figure 4(a) shows the relative win rates and Elo ratings. Interestingly, different from the result in Section 5.2, Minitron-8B consistently outperforms all other LLMs and shows the best Elo rating. This may imply that when multiple agents are involved in the environment, adversarial actions can change the game dynamics. For StarCraft II, we conduct pairwise battles among 7 LLMs, with each pair competing in a single round. Both agents were assigned the same race, Protoss, in all matches. As in Figure 4(b), Claude-3.7-Sonnet shows the best Elo rating, while GPT-4o and Minitron-8B form the second group.
@@ 5.4 Ablation Study for Agentic Modules
[…]
@@ 5.5 Effect of Visual Input
Setup.
To study the effect of visual input, we divide games into two groups: Group 1 (Table 6) includes games where the provided textual game state can entirely be derived from a visual screenshot. Group 2 (Table 6) comprises games where the textual state contains information beyond what is visible in the current frame (e.g., abstracted inventory lists or off-screen character/item details); for this group, we exclude evaluation on Image-only input types for fairness. Minecraft was excluded since the environment does not provide the visionary view of a bot that LLM plays.
Result.
As shown in Table 6, relying solely on Image-only input led to a substantial drop in performance. This was consistently reflected across all models, with average ranks deteriorating significantly compared to Text-only input. In contrast, as shown in Table 66, utilizing Both text and visual inputs produced mixed effects on game scores and model ranks. For instance, in games like Street Fighter III (Group 1), since on-screen visual details are challenging to fully convey textually, adding visual context significantly benefited Claude, increasing its score by $16.6$. Conversely, in narrative-heavy games such as Ace Attorney (Group 2), the same approach often proved detrimental; GPT-4o’s score, for example, dropped by $31.8$, and its average rank in that group fell from 2.9 (Text-only) to 5.0. This highlights that the impact of combining modalities varies considerably, with some scenarios showing improved ranks or scores while others demonstrated a decline.
@@
[…]
Table 5: Modality Comparison (Group 1).
@@
[…]
Table 6: Modality Comparison (Group 2).